# Learning to Translate Noise for Robust Image Denoising

### v1

@@ 3.2.3 Gaussian Injection Block
As illustrated in Figure 2, our noise translation network is built upon a lightweight U-Net architecture, where each layer is composed of Gaussian Injection Blocks (GIBlock). GIBlock incorporates a Nonlinear Activation-Free (NAF) block from NAFNet [6] along with our key idea to align the discrepancy between training and inference stage: Gaussian noise injection.
In the training stage of noise translation network, random levels of Gaussian noise is augmented to an input image $I$ to address diverse noise conditions, enhancing the robustness of the noise translation network. In contrast, in inference stage, noisy input images are given directly to the noise translation network without adding extra Gaussian noise, because the direct noise augmentation to the input images degrades output quality. To establish a consistent Gaussian noise prior in both the training and inference stage, we inject Gaussian noise into every intermediate block of the noise translation network. Since the noise translation network is designed based on U-Net with residual connections between the input $I$ and the output $I_{\mathcal{T}}$, the distortion of the signal caused by injected Gaussian noise is alleviated, while allowing the noise translation network to utilize the Gaussian prior for transforming unseen real noise. Our ablation studies further demonstrate that the proposed Gaussian noise injection is crucial for the noise translation network to effectively translate unseen noise into Gaussian noise during the inference stage.
@@
Denoising Method Metric SIDD Poly CC HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
R2R* [27] PSNR 35.06 36.81 35.26 37.33 39.19 38.29 39.32 41.46 35.36 37.88
SSIM 0.9150 0.9722 0.9756 0.9712 0.9606 0.9663 0.9739 0.9729 0.9664 0.9699
AP-BSN* [15] PSNR 36.32 35.88 33.13 36.66 39.82 37.01 39.04 40.04 33.37 36.87
SSIM 0.9281 0.9751 0.9732 0.9777 0.9766 0.9628 0.9746 0.9798 0.9548 0.9718
SSID* [17] PSNR 37.39 37.13 34.93 38.24 40.85 37.22 39.10 42.89 34.25 38.08
SSIM 0.9338 0.9799 0.9805 0.9808 0.9816 0.9652 0.9738 0.9878 0.9562 0.9757
Restormer* [36] PSNR 39.93 37.63 36.31 38.24 40.05 38.36 39.49 44.02 35.62 38.85
SSIM 0.9598 0.9790 0.9805 0.9753 0.9727 0.9671 0.9768 0.9889 0.9706 0.9745
Restormer-ours PSNR 39.08 38.74 37.60 40.06 41.62 39.68 40.55 44.12 36.14 39.81
SSIM 0.9558 0.9846 0.9861 0.9851 0.9751 0.9761 0.9794 0.9849 0.9747 0.9807
NAFNet* [6] PSNR 39.88 37.13 35.81 38.27 40.16 37.68 39.57 43.48 34.97 38.38
SSIM 0.9594 0.9714 0.9809 0.9784 0.9700 0.9676 0.9781 0.9824 0.9681 0.9746
NAFNet-ours PSNR 39.17 38.67 37.82 39.94 41.94 39.74 40.45 44.17 36.14 39.86
SSIM 0.9566 0.9851 0.9876 0.9853 0.9805 0.9778 0.9796 0.9869 0.9745 0.9822
KBNet* [42] PSNR 40.26 36.79 35.21 38.05 37.93 35.14 37.73 41.65 34.23 37.09
SSIM 0.9618 0.9785 0.9808 0.9785 0.9526 0.9459 0.9657 0.9784 0.9640 0.9681
KBNet-ours PSNR 39.06 38.57 37.59 39.83 41.63 39.71 40.46 44.04 36.04 39.73
SSIM 0.9559 0.9840 0.9859 0.9845 0.9752 0.9773 0.9794 0.9839 0.9739 0.9805
XFormer* [38] PSNR 39.89 37.41 35.93 37.82 39.92 38.30 39.54 43.74 35.52 38.52
SSIM 0.9599 0.9774 0.9781 0.9738 0.9725 0.9673 0.9773 0.9886 0.9702 0.9756
XFormer-ours PSNR 38.98 38.67 37.73 40.22 42.04 39.79 40.61 44.26 36.23 39.94
SSIM 0.9549 0.9852 0.9857 0.9857 0.9816 0.9773 0.9805 0.9879 0.9748 0.9824
@@ 4.1 Experimental Settings
Since our approach is model-agnostic, we employ existing architectures such as NAFNet [6], Restormer [36], KBNet [42] and Xformer [38] as our image denoising network, which is pretrained on BSD400 [24], WED [21], and SIDD medium [1] datasets. BSD400 and WED datasets consist of clean images, while SIDD dataset is composed of real noisy-clean image pairs. During the training of a denoising network, noisy images are created by adding Gaussian noise with a standard deviation of 15 to clean images from the BSD400 and WED datasets, or to different formats of noisy images from the SIDD dataset. Each training batch consists of images drawn equally from two sources: half from the BSD400 and WED datasets, and the other half from the SIDD dataset. The denoising models are trained for 200K iterations with a batch size of 32, except for Restormer and Xformer, where the batch size is reduced to 4 due to the limitation of computational resources. The noise translation network is trained subsequently, with the noisy images from the SIDD dataset augmented by adding stochastic level of Gaussian noise with a range of 0 to 15. The noise translation network is trained for 5K iterations with a batch size of 4. Denoising network and noise translation network are trained with the AdamW [20] optimizer with an initial learning rate of $10^{-3}$, which is reduced to $10^{-7}$ and $10^{-5}$ by a cosine annealing schedule, respectively. Each image is randomly cropped to $256\times 256$ for training. The hyperparameters in our experiment are set as follows: a Gaussian noise injection level $\tilde{\sigma}=100$, $\alpha=5\times 10^{-2}$, and $\beta=2\times 10^{-3}$. All trainings were conducted using two NVIDIA RTX A6000 GPUs.
Refer to caption Poly Noisy Refer to caption AP-BSN 42.04 dB Refer to caption SSID 42.05 dB Refer to caption Restormer 40.72 dB Refer to caption Ground Truth
Refer to caption NAFNet 39.97 dB Refer to caption KBNet 40.45 dB Refer to caption Xformer 40.29 dB Refer to caption Ours 43.59 dB
Refer to caption CC Noisy Refer to caption AP-BSN 35.41 dB Refer to caption SSID 38.45 dB Refer to caption Restormer 37.40 dB Refer to caption Ground Truth
Refer to caption NAFNet 36.99 dB Refer to caption KBNet 38.28 dB Refer to caption Xformer 36.82 dB Refer to caption Ours 42.18 dB
Refer to caption HighISO Noisy Refer to caption AP-BSN 28.25 dB Refer to caption SSID 35.57 dB Refer to caption Restormer 37.79 dB Refer to caption Ground Truth
Refer to caption NAFNet 37.36 dB Refer to caption KBNet 36.96 dB Refer to caption Xformer 37.11 dB Refer to caption Ours 40.37 dB
Figure 3: Comparison between the qualitative results of various denoising networks including ours (noise translation network with pretrained NAFNet), on the out-of-distribution (OOD) datasets. Our result displays cleaner outputs compared to other state-of-the-art networks trained on a real-noise dataset with self-supervised and supervised methods.
@@
To evaluate the generalization performance of our framework, we employ various real-world image denoising benchmarks. We conduct experiments using SIDD validation dataset [1], Poly [31], CC [26], HighISO [33], iPhone, Huawei, Oppo, Sony, and Xiaomi [13]. The SIDD validation dataset consists of images with a resolution of 256$\times$256 pixels, while the Poly and CC datasets contain images with a resolution of $512\times 512$ pixels. Images from HighISO, iPhone, Huawei, Oppo, Sony, and Xiaomi are $1024\times 1024$ pixels in size.
[…]
@@
Input Metric SIDD   Poly    CC HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
$I$ PSNR 37.77 15.24 33.76 21.18 40.13 8.68 8.45 6.35 9.33 17.89
@@
$I_{\mathcal{T}}$ PSNR 39.17 38.67 37.82 39.94 41.94 39.74 40.45 44.17 36.14 39.86
SSIM 0.9566 0.9851 0.9876 0.9853 0.9805 0.9778 0.9796 0.9869 0.9745 0.9822
4.2 Results
Denoising performance on real noise
Table 1 presents the performance of the proposed approach applied to the denoising networks, including Restormer [36], NAFNet [6], KBNet [42], and Xformer [38], alongside results from self-supervised image denoising methods, including R2R [27], AP-BSN [15], and SSID [17] for comparison. For evaluating existing methods, we use the officially published models trained on the SIDD dataset, where self-supervised methods are trained with only noisy images. As shown in Table 1, incorporating our noise translation framework—denoted by Restormer-ours, NAFNet-ours, KBNet-ours, and Xformer-ours—yields significant average PSNR improvements of 0.96dB, 1.48dB, 2.64dB, and 1.42dB, respectively, in out-of-distribution (OOD) scenarios. This demonstrates the effectiveness of our method in enhancing the generalization performance of real-world image denoising networks. Additional comparisons with recent studies on generalizable denoising, including LAN [12] and AFM [29], are presented in the Appendix.
Qualitative results on OOD benchmarks
Figure 3 illustrates qualitative results for denoising models on several real-world OOD datasets, including Poly [31], CC [26], and HighISO [33]. Our model demonstrates significant superiority over other denoising methods in these challenging scenarios. Additional qualitative results on every benchmark listed in Table 1 are available in the Appendix.
4.3 Analysis
We perform a comprehensive analysis using the noise translation network with pretrained NAFNet (NAFNet-ours).
@@ 4.3 Analysis
We assess the effectiveness of our noise translation network by comparing it to simply adding Gaussian noise to the input. As shown in table 2, adding Gaussian noise to the input achieve reasonable generalization performance. However, some datasets perform better with an input of $I+\mathcal{N}_{10}$, while others perform better with $I+\mathcal{N}_{15}$. This variation suggests that each image or dataset exhibits distinct noise characteristics, highlighting the need for a more adaptable approach than applying a fixed level of Gaussian noise. Our noise translation network flexibly transforms the input noise of each image into ideal Gaussian noise, resulting in significant performance improvements across all datasets.
Table 3: Effects of using Gaussian injection block and explicit noise translation loss. The results for each OOD dataset are detailed in the Appendix.
Metric SIDD OOD Avg.
Baseline Translation PSNR 39.35 39.27
SSIM 0.9573 0.9774
+ Gaussian Injection Block PSNR 39.05 39.61
SSIM 0.9556 0.9801
+ Explicit Noise Translation Loss PSNR 39.37 38.66
SSIM 0.9574 0.9720
+ Both (Ours) PSNR 39.17 39.86
SSIM 0.9566 0.9822
Gaussian injection and explicit noise translation loss
Table 3 illustrates the performance gains attributed to each component of the proposed method. The baseline translation in Table 3 refers to the results obtained by applying our method without Gaussian noise injection and explicit noise translation loss. We observe the best performance gains when both Gaussian noise injection and explicit noise translation loss are both applied.
Table 4: Quantitative results to demonstrate the adaptability of our noise translation framework. The metrics for each OOD dataset are presented in the Appendix.
$\mathcal{D}_{\theta^{*}}-train$ $\mathcal{D}_{\theta^{*}}-test$ Metric SIDD OOD Avg.
NAFNet [6] Restormer [36] PSNR 39.19 39.75
SSIM 0.9567 0.9820
KBNet [42] PSNR 39.15 39.84
SSIM 0.9565 0.9819
Xformer [38] PSNR 39.02 39.84
SSIM 0.9555 0.9822
Adaptability of the noise translation network
In our framework, the noise translation network and the denoising network perform distinct tasks, enabling the use of a single noise translation network with different pretrained denoising networks. Table 4 demonstrates the flexibility by applying a noise translation network trained with pretrained NAFNet [6] to other architectures: Restormer [36], KBNet [42], and Xformer [38]. The results show comparable performance to using a noise translation network specifically trained for each denoising model, reported in Table 1. This suggests that a single-time trained noise translation network can be employed with various pretrained denoising networks, without a necessary additional training.
Analysis of translated noise
Figure 4 visualizes the results of our noise translation process. Real noise exhibits strong spatial correlation and signal dependency. Our noise translation network effectively alleviates these characteristics, converting the noise into a form that resembles ideal Gaussian noise. Figure 5 presents the analysis of the noise distribution of two images dealt in Figure 4. In the spatial domain, ideal Gaussian noise follow a normal distribution, while in the frequency domain, it follows Rayleigh distribution, as discussed in Section 3.2.2. The original real-noise distribution significantly deviates from the expected target. After the translation, the noise closely matches the target distributions, demonstrating the effectiveness of our method in aligning the noise with spatially uncorrelated, i.i.d Gaussian noise.
@@
Figure 4: Visual results of noise translation. The noisy image in the top row is from the Poly [31] dataset, while the one in the bottom row is from the CC [26] dataset. For enhanced visualization, noise is displayed as the absolute value, scaled by a factor of 10.
@@
Evaluating Trade-offs in ID Performance
Figure 6 displays an enlarged qualitative results on the SIDD validation dataset. Although our method may seem to show a trade-off in in-distribution (ID) performance, this is primarily attributed to the pronounced overfitting of other models, which often reconstruct irrelevant artifacts inherent in the training data. In contrast, our approach preserves high visual quality while avoiding overfitting, effectively removing noise without introducing such artifacts.
Table 5: Parameter counts and MACs of denoising networks and our noise translation network. We report the number of parameters and MACs at inference, estimated with an input size of 256$\times$256.
@@
Computational efficiency
Table 5 presents a comparison of our noise translation network with other image denoising networks in terms of the number of parameters and multiply–accumulate operations (MACs). Our noise translation network is significantly smaller in size compared to the image denoising networks, resulting in negligible additional computational cost during inference.

### v2

@@ 3.2.3 Gaussian Injection Block
As shown in Figure 3, our noise translation network is based on a lightweight U-Net architecture, where each layer contains a Gaussian Injection Block (GIBlock). Each GIBlock combines a Nonlinear Activation-Free (NAF) block [6] with our key design, the injection of Gaussian noise, to consistently impose a Gaussian prior throughout the network.
While injecting Gaussian noise into the input can provide a Gaussian prior to some extent, it inevitably distorts the input signal. In contrast, our method injects Gaussian noise internally into sub-blocks of the network, allowing the network to receive the Gaussian prior gradually without perturbing the original image for translating unseen real noise effectively. Additionally, the residual connections between the input $I$ and the output $I_{\mathcal{T}}$ further mitigate any signal distortion introduced by the injected noise. Ablation studies confirm that Gaussian noise injection is essential for enabling the noise translation network to reliably map unseen noise to a Gaussian distribution during inference.
@@
Denoising Algo. Metric SIDD Poly $C$ HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
Self-supervised R2R* [32] PSNR 35.09 36.84 35.28 37.37 39.25 38.35 39.38 41.55 35.39 37.93
SSIM 0.9154 0.9726 0.9758 0.9716 0.9614 0.9667 0.9744 0.9735 0.9668 0.9703
AP-BSN* [19] PSNR 36.35 35.91 33.15 36.70 39.90 37.06 39.11 40.13 33.39 36.92
SSIM 0.9285 0.9755 0.9734 0.9781 0.9773 0.9633 0.9750 0.9805 0.9552 0.9723
SSID* [21] PSNR 37.43 37.17 34.95 38.28 40.95 37.27 39.17 43.03 34.27 38.14
SSIM 0.9343 0.9803 0.9806 0.9811 0.9823 0.9656 0.9743 0.9883 0.9566 0.9762
APR-RD* [16] PSNR 38.06 37.01 35.83 38.75 40.40 36.75 38.93 43.35 33.74 38.10
SSIM 0.9470 0.9781 0.9824 0.9823 0.9763 0.9590 0.9713 0.9861 0.9483 0.9730
[…]
SSIM 0.9529 0.9808 0.9810 0.9793 0.9801 0.9694 0.9790 0.9899 0.9697 0.9786
[…]
SSIM 0.9446 0.9795 0.9819 0.9782 0.9730 0.9702 0.9789 0.9857 0.9690 0.9771
AFM* [34] PSNR 38.29 37.71 36.81 39.12 40.56 38.33 40.13 44.66 35.25 39.07
SSIM 0.9474 0.9800 0.9828 0.9797 0.9769 0.9679 0.9795 0.9901 0.9665 0.9745
[…]
SSIM 0.9354 0.9764 0.9754 0.9764 0.9771 0.9593 0.9754 0.9885 0.9563 0.9731
Ours NAFNet* [6] PSNR 39.97 37.17 35.69 38.32 40.25 37.73 39.64 43.65 34.99 38.43
SSIM 0.9600 0.9717 0.9811 0.9788 0.9707 0.9680 0.9786 0.9829 0.9685 0.9750
NAFNet + NTN PSNR 39.24 38.72 37.84 40.00 42.08 39.83 40.55 44.34 36.17 39.94
SSIM 0.9570 0.9855 0.9877 0.9856 0.9812 0.9782 0.9801 0.9875 0.9749 0.9826
XFormer* [45] PSNR 39.98 37.45 35.95 37.86 39.99 38.35 39.61 43.89 35.54 38.58
SSIM 0.9603 0.9778 0.9782 0.9741 0.9732 0.9678 0.9779 0.9892 0.9706 0.9761
XFormer + NTN PSNR 39.10 38.75 37.82 40.29 42.20 39.89 40.70 44.44 36.25 40.04
SSIM 0.9557 0.9856 0.9862 0.9860 0.9826 0.9779 0.9807 0.9886 0.9751 0.9828
@@ 4.1 Experimental Settings
CC Noisy
R2R 37.03dB
AP-BSN 35.43dB
SSID 38.48dB
APR-RD 39.77dB
Mask-DN 38.05dB
CLIP-DN 37.87dB
AFM 39.28dB
IDF 35.80dB
NAFNet 37.98dB
XFormer 36.85dB
Ours 42.19dB
Ground Truth
HighISO Noisy
R2R 34.52dB
AP-BSN 28.26dB
SSID 35.59dB
APR-RD 35.70dB
Mask-DN 36.83dB
CLIP-DN 36.97dB
AFM 38.30dB
IDF 33.80dB
NAFNet 37.39dB
XFormer 37.14dB
Ours 40.43dB
Ground Truth
Figure 4: Comparison between the qualitative results of various denoising networks including ours (noise translation network with pretrained NAFNet), on the out-of-distribution (OOD) datasets. Our result displays cleaner outputs compared to other state-of-the-art networks that are directly trained from a real-noise dataset. Zoom in for better comparison.
Our framework is model-agnostic and can be applied to various denoising networks. To demonstrate the effectiveness of the proposed approach, we adopt NAFNet [6] and Xformer [45], pretrained on BSD400 [29], WED [25], and SIDD-Medium [1] datasets. BSD400 and WED consist of clean images while SIDD provides real noisy-clean image pairs. During the training of a denoising network, noisy images are created by adding Gaussian noise with a standard deviation of 15 to clean images from BSD400 and WED and noisy images from SIDD. The noise translation network is trained subsequently, using SIDD images with data augmentation of random Gaussian noise sampled from $[0,15]$. The hyperparameters are set as follows: a Gaussian noise level $\tilde{\sigma}=100$, $\alpha=5\times 10^{-2}$, and $\beta=2\times 10^{-3}$. Additional training details such as batch size or training resources are discussed in Appendix.
@@
To evaluate the generalization performance of our framework, we employ various real-world image denoising benchmarks. We conduct experiments using SIDD validation dataset [1], Poly [38], CC [31], HighISO [40], iPhone, Huawei, Oppo, Sony, and Xiaomi [17]. The SIDD validation dataset consists of $\times256$ images, while the Poly and CC datasets contain images with a resolution of $512\times 512$ pixels. Images from HighISO, iPhone, Huawei, Oppo, Sony, and Xiaomi are $1024\times 1024$ pixels in size.
4.2 Results
Performance on real noise
Table 1 presents the performance of the proposed approach applied to existing denoising networks, NAFNet [6] and Xformer [45]. We also compare their results with those from self-supervised image denoising methods, including R2R [32], AP-BSN [19], SSID [21] and APR-RD [16] and generalizable image denoising methods, including Mask-Denoising [5], Clip-Denoising [8], AFM [34], and IDF [15]. We used the officially published, SIDD-pretrained models for comparisons if available. For the methods without such pretrained models such as Mask-Denoising, Clip-Denoising, and IDF, we trained these models on the SIDD training dataset until convergence to ensure a fair comparison.
As shown in Table 1, incorporating our noise translation framework, denoted by NAFNet + NTN and Xformer + NTN, achieves substantial average PSNR gains of 1.51dB and 1.46dB, respectively, in out-of-distribution (OOD) scenarios. Furthermore, our method demonstrates superior real-world image denoising generalization performance compared to other self-supervised and generalizable denoising methods. Additional comparisons—including LAN [14] applied to the same NAFNet architecture, as well as results on Restormer [43] and KBNet [49] with NTN—are provided in the Appendix.
Qualitative results on OOD benchmarks
Figure 4 illustrates qualitative results for denoising models on several real-world OOD datasets, including CC [31], and HighISO [40]. Our model demonstrates significant superiority over other denoising methods in these challenging scenarios. Qualitative results on every benchmark listed in Table 1 are available in the Appendix.
4.3 Analysis
In this section, we present a comprehensive analysis using our noise translation network combined with a pretrained NAFNet as a denoising network (NAFNet + NTN).
Table 2: Ablation study of the Gaussian injection block (GIBlock) and the explicit noise translation loss. Full results on individual OOD datasets are provided in the Appendix.
Metric SIDD OOD Avg.
Baseline translation PSNR 39.35 39.27
SSIM 0.9573 0.9774
+ GIBlock PSNR 39.05 39.61
SSIM 0.9556 0.9801
+ Explicit loss ($\mathcal{L}_{\text{explicit}}$) PSNR 39.33 39.61
SSIM 0.9572 0.9809
+ Both (Ours) PSNR 39.24 39.94
SSIM 0.9570 0.9826
Table 3: Quantitative results based on variations in the noisy input to the pretrained denoising network on the in-distribution SIDD validation set and other out-of-distribution real-world benchmarks. II, $I+\mathcal{N}_{5}$, $I+\mathcal{N}_{10}$, and $I+\mathcal{N}_{15}$ represent the noisy input images with additional Gaussian noise levels of 0, 5, 10, and 15, respectively, which are fed into the pretrained Gaussian denoising network.
@@ 4.3 Analysis
Input Metric SIDD   Poly    $C$ HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
II PSNR 37.77 15.24 33.76 21.18 40.13 8.68 8.45 6.35 9.33 17.89
@@
$I_{\mathcal{T}}$ PSNR 39.24 38.72 37.84 40.00 42.08 39.83 40.55 44.34 36.17 39.94
SSIM 0.9570 0.9855 0.9877 0.9856 0.9812 0.9782 0.9801 0.9875 0.9749 0.9826
Ablation study
Table 2 presents the ablation study results of the proposed algorithm. The baseline model is trained using only the implicit noise translation loss, excluding both the Gaussian noise injection block (GIBlock) and the explicit noise translation loss. The results demonstrate that both components (GIBlock and explicit loss) are essential for enhancing performance: GIBlock promotes the translation of complex noise into Gaussian noise, while the explicit loss provides a clear constraint that guides the generation toward i.i.d. Gaussian noise.
@@
We evaluate the effectiveness of our noise translation network by comparing it to directly adding Gaussian noise to the input. As shown in Table 3, simply adding Gaussian noise yields reasonable generalization. However, some datasets perform better with $I+\mathcal{N}_{10}$, while others with $I+\mathcal{N}_{15}$, indicating that noise characteristics vary across datasets or even individual images. This variation underscores the limitation of fixed noise and the need for a more adaptive approach. Our noise translation network can adaptively transform input noise into ideal Gaussian noise, achieving consistent performance gains across all datasets.
[…]
$\mathcal{D}_{\theta^{*}}$ for testing $\mathcal{D}_{\theta^{*}}$ for training $\mathcal{T}_{\phi}$ SIDD OOD Avg.
Xformer [45] NAFNet [6] 39.17 39.94
Xformer [45] 39.10 40.04
Transferability of noise translation network
The noise translation network is functionally decoupled from the denoising network, allowing a single trained noise translation model to be paired with various pretrained denoising networks. Table 4 demonstrates this property by applying the noise translation network, trained with pretrained NAFNet [6], to other architecture, Xformer [45]. The performance remains comparable to that of individually trained noise translation networks for each denoiser. This suggests that a one-time trained noise translation network can generalize effectively across diverse pretrained denoising networks without additional training.
@@
Figure 5: Visual results of noise translation. The noisy image in the top row is from the Poly [38] dataset, while the one in the bottom row is from the $C$ [31] dataset. For enhanced visualization, noise is displayed as the absolute value, scaled by a factor of 10.
@@
(a) $C$
(b) HighISO
Figure 7: Analysis of noise characteristics. The blue curve illustrates the changes in noise level and auto-correlation resulting from adding various levels of Gaussian noise to the real noise. The results are averaged over the $C$ [31] and HighISO [40] datasets.
Analysis of translated noise
Figure 5 visualizes the effect of our noise translation. Real noise shows strong spatial correlation, which our network effectively removes, producing noise that closely resembles ideal Gaussian noise.
Figure 6 shows the noise distributions of the two images in Figure 5. In the spatial domain, ideal Gaussian noise follows a normal distribution; in the frequency domain, it follows a Rayleigh distribution (see Section 3.2.2). Real noise significantly deviates from these targets, whereas translated noise closely matches them, demonstrating effective alignment with spatially uncorrelated, i.i.d. Gaussian noise.
Additionally, Figure 7 plots the noise level versus auto-correlation for each dataset, comparing three cases: (1) real noise, (2) real noise with added Gaussian noise, and (3) translated noise produced by our method. While adding Gaussian noise to real noise increases the overall noise level and reduces auto-correlation, our translated noise achieves substantially lower auto-correlation at comparable noise levels. This indicates that our translation process more effectively decorrelates the noise.
Evaluating Trade-offs in ID performance
Figure 8 displays enlarged qualitative results on the SIDD validation dataset. Although our method may appear to exhibit a trade-off in in-distribution performance, this primarily stems from the pronounced overfitting of other models, which often reconstruct irrelevant artifacts inherent in the training data. In contrast, our approach avoids such overfitting while still producing high-quality reconstructions.
If improving in-distribution performance on a specific dataset is desired, the denoising network can be fine-tuned. We demonstrate this by freezing the trained translation network and fine-tuning only the denoising network on the SIDD dataset. As shown in Table 5, this procedure yields an improvement of approximately 0.3dB on SIDD while having negligible impact on out-of-distribution performance.
Computational efficiency
Table 6 presents a comparison of our noise translation network with other image denoising networks including NAFNet [6] and Xformer [45] in terms of the number of parameters and multiply-accumulate operations (MACs). Our noise translation network is substantially smaller in both parameter count and MACs than the image denoising networks, adding only negligible computational overhead during inference.
@@
Table 5: Impact of SIDD [1] fine-tuning. We present the performance of NAFNet and Xformer when the noise translation network is frozen and only the denoising network is fine-tuned.
Metric SIDD OOD Avg.
NAFNet + NTNet PSNR 39.56 (+0.32) 39.95 (+0.01)
SSIM 0.9580 (+0.001) 0.9830 (+0.0004)
Xformer + NTNet PSNR 39.43 (+0.33) 39.97 (-0.07)
SSIM 0.9589 (+0.0032) 0.9903 (+0.0075)
Table 6: Parameter counts and MACs of denoising networks and our noise translation network. We report the number of parameters and MACs at inference, estimated with an input size of $\times256$.